A silver pin is exactly 4.00 cm long when its temperature is 177∘C Find the decrease in long of the pin when it cools to 26.0∘C
Answer #1

dL = L*alpha*dT

where alpha = thermal coefficient of silver = 18*10^-6 / oC

=> dL = 0.04*18*10^-6*(26 - 177)

=> dL = - 108.72*10^-6 m = - 0.0108 cm  

- ve sign shows decrease.